OpenStudy (avrieltkd):

What is the simplified expression for 4 to the power of negative 3 multiplied by 3 to the power of 4 multiplied by 4 to the power of 2 whole over 3 to the power of 5 multiplied by 4 to the power of negative 2 ?

OpenStudy (hoot_hoot):

\[\frac{ (\frac{ 1 }{ 4^{3} })(3^{4})(4^{2}) }{(3^{5})(\frac{ 1 }{ 4^{2} }) }\] Simplify. The numerator will be: \[\frac{ 3^{4} }{ 4 }\] Find their equivalent values. For example: \[2^{2} = 4\] Then, follow the operation.

OpenStudy (mathmale):

I'd suggest you avoid presenting your math expression in words unless specifically told to do so. Instead, use mathematical symbols:\[\frac{ (\frac{ 1 }{ 4^{3} })(3^{4})(4^{2}) }{(3^{5})(\frac{ 1 }{ 4^{2} }) }\] (as presented by hoot_hoot).

OpenStudy (mathmale):

Start by reducing\[\frac{ 3^4 }{ 3^5 }\]